# Count Every Rotation and Every Rotation Counts: Exploring Drone Dynamics via Propeller Sensing

### v1

@@ 5.1. Implementation
As shown in Fig.14, we implement EventPro with a commercial event camera, the Prophesee EVK4 HD evaluation kit (Prophesee, 2024), which is equipped with the IMX636ES event-based vision sensor for high-definition event data capture (1280 × 720 pixels) using the Soyo SFA0820-5M lens. Our implementation contains more than 20000 LOC (lines of Python code). EventPro operates on a laptop equipped with an Intel i7-12900K CPU, and 32GB RAM, running on the Ubuntu 20.04 operating system. The event camera is connected to the laptop via USB 3.0. All data sampling, streaming, and computation are performed in real-time on a laptop, enabling accurate and efficient speed estimation of the drone’s propellers.
@@ 5.2.1. Experimental Methodology
$\bullet$ Outdoor experiments. We adopt a commercial delivery drone and conduct real-world experiments in a test field. This drone is manufactured by a company specializing in daily and city-wide instant delivery services. The drone features six propellers attached to brushless motors and is controlled by a PX4(PX4, 2024) flight controller. An event camera was positioned on a ground-based bracket to capture the drone in flight. We conducted extensive experiments over 10 hours, collecting more than 600GB of raw event data.
@@
The ground truth for indoor experiments is obtained using a high-precision laser tachometer (UNI-T UT372(UNI-T, 2024)) (Fig.14a) with a relative error of ±0.04%. A reflective marker is affixed to the motor shell to reflect infrared light, and the tachometer is mounted on a tripod aimed at the marker. For outdoor experiments, a built-in magnetic encoder attached to the shaft of the drone’s motor provides ground truth with a relative error of ±0.5%. All measurement are synchronized with EventPro and baselines for comparison.
Baselines. We compare EventPro with two approaches: 1) Ev-Tach (Zhao et al., 2023), a state-of-the-art system that leverages the classicial iterative closest point (ICP) algorithm for rotational speed estimation; 2) EEPPR (Kolář et al., 2025), a cutting-edge method that measures the period of rotational motion by analyzing the 3D correlation of events within a spatiotemporal window.
@@ 5.2.3. System Robustness Evaluation
Impact of rotational speed. Fig.16a depicts the impact of rotational speed within the typical operating range of drone (Deters et al., 2017). EventPro consistently outperforms baselines across all tested settings. As rotational speed increases, we observe a slight downward trend across all methods. This is due to slower speeds amplifying the impact of the same error, leading to higher RMAE. In contrast, EventPro consistently maintains stable and satisfactory estimation performance regardless of rotational speed changes.
@@ 5.4.1. Experimental Methodology
The experiment setup is depicted in Fig.14b. A Real-Time Kinematic Positioning (RTK) base station was deployed to provide high-fidelity localization results as the ground truth. Following standard practice in precision landing systems (Wang et al., 2025c), we adopt mean localization error in meters as the primary metric, defined as the average positioning error across latitude, longitude, and altitude. This metric directly reflects landing performance: larger errors may result in landing pad misalignment, failed connection with charging ports, or necessitate abort maneuvers. We compare the mean error of EventPro against three baselines: ($i$) GPS; ($ii$) MicNest (Wang et al., 2022), a state-of-the-art acoustic localization method; and (iiiiii) mmE-loc (Wang et al., 2025c) a state-of-the-art localization system based on event cameras and mmWave radar fusion.
Flight trajectories. The drone operates automatically according to pre-planned trajectories from the ground station. We consider two classical flight trajectories following previous work (Wang et al., 2022): 1) Vertical flight, where the drone takes off, climbs vertically to a specified altitude, and then lands back onto the platform; 2) Spiral flight, where the drone ascends to a fixed altitude, performs three loops along a complex squared spiral path, and returns to land on the platform. The default flight speed is 4m/sm/s.
@@ 6. Related Work
Rotational speed estimation. Rotational motion sensing methods are broadly categorized as invasive or non-invasive. Invasive methods, such as mechanical tachometers (Liptak, 2003), electrostatic sensors (Li et al., 2019), Hall-effect sensors (Kathirvelan et al., 2014), and optical encoders (Asfia and Rezaei, 2011), involve physical contact or additional hardware, potentially disrupting the system’s natural state and limiting flexibility. Non-invasive methods utilize acoustic (Li et al., 2018), RF signals (Heggo et al., 2022), WiFi (Lin et al., 2022), optical (Wang et al., 2017, 2018; Kim et al., 2016; Natili et al., 2020; Technology, 2022), and radar (Ciattaglia et al., 2023b, a; Ding et al., 2023) techniques, but they face significant limitations in spatial sensing range or temporal resolution, often struggling with noise interference. For example, laser tachometers (Technology, 2022; Cheng et al., 2011) require retroreflective markers and precise aiming, with an operation distance under 20 cm. RGB camera-based methods (Wang et al., 2017, 2018) are restricted by a 60 fps frame rate, capping measurable speeds at 1500 RPM. Additionally, most of the above methods are designed for scenarios involving a single rotating object, which is inadequate for our settings.
@@
The most relevant work, Ev-Tach (Zhao et al., 2023), introduces a notable rotational speed estimation solution using event cameras. It employs a 3D points-based registration method to estimate the transformation between two slices of event streams, thereby determining the angle of rotation. However, Ev-Tach primarily targets fixed rotating objects and requires a small-scale work environment with a maximum sensing distance of around 70 cm. In contrast, our EventPro focuses on propeller speed sensing for high-speed multirotors in real-world environments. This shift from fixed targets at close distances to dynamic targets in diverse conditions, combined with outdoor noise, introduces new challenges to fully harness event cameras’ potential, as elaborated in $\S$1.
@@
Event-based algorithms and systems. Event camera-based sensing has garnered increasing popularity across various applications, owing to their high temporal resolution, low latency, and high dynamic range. Recent systems employ them for object localization (Xu et al., 2023a; Wang et al., 2025a; Luo et al., 2024), object tracking (Hao et al., 2023; Falanga et al., 2020; Mitrokhin et al., 2018), SLAM (Rebecq et al., 2016), and motion estimation (Gallego and Scaramuzza, 2017; Ng et al., 2023; Kim and Kim, 2021; Shiba et al., 2024; Zhu et al., 2019). Note that though (Gallego and Scaramuzza, 2017; Ng et al., 2023; Kim and Kim, 2021) focus on rotational speed estimation, they aim to measure the rotation of the event camera itself. Our work fundamentally differs by leveraging the event camera to estimate the rotational speed of external targets in real-world scenarios. In EventPro, we introduce a novel motion compensation-based pipeline that incorporates propeller geometries for accurate and real-time drone propeller speed estimation, featuring three effective modules. Our design aims to improve the accuracy and efficiency of event-based systems, particularly benefiting high-speed drones.
@@ 7. Discussion
Deployment Conditions and Future Applications. EventPro is applicable to various multi-rotor platforms due to their similar rotational characteristics. However, event cameras remain in active development with certain limitations, including low resolution and fixed aperture. Based on field experiments, we identify optimal deployment conditions: observation distances under 50m, viewing angles between 0-50 degrees, and illumination below 100,000 lux. These conditions are well-suited for drone landing scenarios (Zhao et al., 2025; Wang et al., 2025d). Looking forward, EventPro holds potential for broader applications, including multi-drone localization (Chen et al., 2024) and high-speed FPV trajectory prediction. To enable these applications, the system requires enhancement with clustering algorithms for distinguishing multiple drones and more robust spatiotemporal feature extraction methods for detecting small-scale drones at extended distances.
Influence of Weather. Illumination is the primary environmental factor affecting event camera performance (Wang et al., 2025b). Extreme lighting conditions on sunny days have minimal impact on EventPro, as event cameras respond to relative rather than absolute light intensity changes. Conversely, weak illumination reduces event generation due to insufficient intensity variations, as demonstrated in Fig.16d. Additionally, dynamic background motion induced by fog or rain may degrade drone localization performance. A future direction is developing more robust event preprocessing algorithms that leverage spatiotemporal drone features in the frequency domain to mitigate these adverse effects (Ruan et al., 2025b, a).
Computational and Cost consideration. Our evaluation deploys EventPro on a laptop equipped with an Intel i7-12900K CPU and 32GB RAM, achieving real-time performance. Notably, modern compact edge computing platforms, such as the ASUS NUC 14 Pro (ASUS, 2025), can fully support these computational requirements, enabling flexible deployment in resource-constrained environments. Regarding cost considerations, the primary bottleneck lies in event camera hardware, with prices ranging from \$300 to \$5,000 depending on resolution, sensitivity. However, we anticipate significant price reductions in the near future, driven by rapid advancements in neuromorphic vision hardware and increasing commercial adoption. Our work presents an active and contactless solution for drone sensing, positioning our system to readily benefit from ongoing advancements in event camera technology.

### v2

@@ 5.1. Implementation
As shown in Fig.14, we implement EventPro with a commercial event camera, the Prophesee EVK4 HD evaluation kit (evk4), which is equipped with the IMX636ES event-based vision sensor for high-definition event data capture (1280 × 720 pixels) using the Soyo SFA0820-5M lens. Our implementation contains more than 20000 LOC (lines of Python code). EventPro operates on a laptop equipped with an Intel i7-12900K CPU, and 32GB RAM, running on the Ubuntu 20.04 operating system. The event camera is connected to the laptop via USB 3.0. All data sampling, streaming, and computation are performed in real-time on a laptop, enabling accurate and efficient speed estimation of the drone’s propellers.
@@ 5.2.1. Experimental Methodology
$\bullet$ Outdoor experiments. We adopt a commercial delivery drone and conduct real-world experiments in a test field. This drone is manufactured by a company specializing in daily and city-wide instant delivery services. The drone features six propellers attached to brushless motors and is controlled by a PX4(px4) flight controller. An event camera was positioned on a ground-based bracket to capture the drone in flight. We conducted extensive experiments over 10 hours, collecting more than 600GB of raw event data.
@@
The ground truth for indoor experiments is obtained using a high-precision laser tachometer (UNI-T UT372(UT372)) (Fig.14a) with a relative error of ±0.04%. A reflective marker is affixed to the motor shell to reflect infrared light, and the tachometer is mounted on a tripod aimed at the marker. For outdoor experiments, a built-in magnetic encoder attached to the shaft of the drone’s motor provides ground truth with a relative error of ±0.5%. All measurement are synchronized with EventPro and baselines for comparison.
Baselines. We compare EventPro with two approaches: 1) Ev-Tach (zhao2023ev), a state-of-the-art system that leverages the classicial iterative closest point (ICP) algorithm for rotational speed estimation; 2) EEPPR (kolavr2024ee3p3d), a cutting-edge method that measures the period of rotational motion by analyzing the 3D correlation of events within a spatiotemporal window.
@@ 5.2.3. System Robustness Evaluation
Impact of rotational speed. Fig.16a depicts the impact of rotational speed within the typical operating range of drone (deters2017static). EventPro consistently outperforms baselines across all tested settings. As rotational speed increases, we observe a slight downward trend across all methods. This is due to slower speeds amplifying the impact of the same error, leading to higher RMAE. In contrast, EventPro consistently maintains stable and satisfactory estimation performance regardless of rotational speed changes.
@@ 5.4.1. Experimental Methodology
The experiment setup is depicted in Fig.14b. A Real-Time Kinematic Positioning (RTK) base station was deployed to provide high-fidelity localization results as the ground truth. Following standard practice in precision landing systems (wang2025ultra), we adopt mean localization error in meters as the primary metric, defined as the average positioning error across latitude, longitude, and altitude. This metric directly reflects landing performance: larger errors may result in landing pad misalignment, failed connection with charging ports, or necessitate abort maneuvers. We compare the mean error of EventPro against three baselines: ($i$) GPS; ($ii$) MicNest (wang2022micnest), a state-of-the-art acoustic localization method; and (iiiiii) mmE-loc (wang2025ultra) a state-of-the-art localization system based on event cameras and mmWave radar fusion.
Flight trajectories. The drone operates automatically according to pre-planned trajectories from the ground station. We consider two classical flight trajectories following previous work (wang2022micnest): 1) Vertical flight, where the drone takes off, climbs vertically to a specified altitude, and then lands back onto the platform; 2) Spiral flight, where the drone ascends to a fixed altitude, performs three loops along a complex squared spiral path, and returns to land on the platform. The default flight speed is 4m/sm/s.
@@ 6. Related Work
Rotational speed estimation. Rotational motion sensing methods are broadly categorized as invasive or non-invasive. Invasive methods, such as mechanical tachometers (liptak2003instrument), electrostatic sensors (li2019digital), Hall-effect sensors (kathirvelan2014hall), and optical encoders (asfia2011high), involve physical contact or additional hardware, potentially disrupting the system’s natural state and limiting flexibility. Non-invasive methods utilize acoustic (li2018pcias), RF signals (heggo2022rftacho), WiFi (lin2022wi), optical (wang2017rotational; wang2018rotational; kim2016visual; natili2020video; lasertacho), and radar (ciattaglia2023uav; ciattaglia2023determination; ding2023rotation) techniques, but they face significant limitations in spatial sensing range or temporal resolution, often struggling with noise interference. For example, laser tachometers (lasertacho; cheng2011contactless) require retroreflective markers and precise aiming, with an operation distance under 20 cm. RGB camera-based methods (wang2017rotational; wang2018rotational) are restricted by a 60 fps frame rate, capping measurable speeds at 1500 RPM. Additionally, most of the above methods are designed for scenarios involving a single rotating object, which is inadequate for our settings.
@@
The most relevant work, Ev-Tach (zhao2023ev), introduces a notable rotational speed estimation solution using event cameras. It employs a 3D points-based registration method to estimate the transformation between two slices of event streams, thereby determining the angle of rotation. However, Ev-Tach primarily targets fixed rotating objects and requires a small-scale work environment with a maximum sensing distance of around 70 cm. In contrast, our EventPro focuses on propeller speed sensing for high-speed multirotors in real-world environments. This shift from fixed targets at close distances to dynamic targets in diverse conditions, combined with outdoor noise, introduces new challenges to fully harness event cameras’ potential, as elaborated in $\S$1.
@@
Event-based algorithms and systems. Event camera-based sensing has garnered increasing popularity across various applications, owing to their high temporal resolution, low latency, and high dynamic range. Recent systems employ them for object localization (xu2023taming; wang2025event; luo2024eventtracker), object tracking (hao2023eventboost; falanga2020dynamic; mitrokhin2018event), SLAM (rebecq2016evo), and motion estimation (gallego2017accurate; ng2023direct; kim2021real; shiba2024secrets; zhu2019unsupervised). Note that though (gallego2017accurate; ng2023direct; kim2021real) focus on rotational speed estimation, they aim to measure the rotation of the event camera itself. Our work fundamentally differs by leveraging the event camera to estimate the rotational speed of external targets in real-world scenarios. In EventPro, we introduce a novel motion compensation-based pipeline that incorporates propeller geometries for accurate and real-time drone propeller speed estimation, featuring three effective modules. Our design aims to improve the accuracy and efficiency of event-based systems, particularly benefiting high-speed drones.
@@ 7. Discussion
Deployment Conditions and Future Applications. EventPro is applicable to various multi-rotor platforms due to their similar rotational characteristics. However, event cameras remain in active development with certain limitations, including low resolution and fixed aperture. Based on field experiments, we identify optimal deployment conditions: observation distances under 50m, viewing angles between 0-50 degrees, and illumination below 100,000 lux. These conditions are well-suited for drone landing scenarios (zhao2025airscape; wang2025mme). Looking forward, EventPro holds potential for broader applications, including multi-drone localization (chen2024soscheduler) and high-speed FPV trajectory prediction. To enable these applications, the system requires enhancement with clustering algorithms for distinguishing multiple drones and more robust spatiotemporal feature extraction methods for detecting small-scale drones at extended distances.
Influence of Weather. Illumination is the primary environmental factor affecting event camera performance (wang2025enabling). Extreme lighting conditions on sunny days have minimal impact on EventPro, as event cameras respond to relative rather than absolute light intensity changes. Conversely, weak illumination reduces event generation due to insufficient intensity variations, as demonstrated in Fig.16d. Additionally, dynamic background motion induced by fog or rain may degrade drone localization performance. A future direction is developing more robust event preprocessing algorithms that leverage spatiotemporal drone features in the frequency domain to mitigate these adverse effects (ruan2025pre; ruan2025edmamba).
Computational and Cost consideration. Our evaluation deploys EventPro on a laptop equipped with an Intel i7-12900K CPU and 32GB RAM, achieving real-time performance. Notably, modern compact edge computing platforms, such as the ASUS NUC 14 Pro (nuc), can fully support these computational requirements, enabling flexible deployment in resource-constrained environments. Regarding cost considerations, the primary bottleneck lies in event camera hardware, with prices ranging from \$300 to \$5,000 depending on resolution, sensitivity. However, we anticipate significant price reductions in the near future, driven by rapid advancements in neuromorphic vision hardware and increasing commercial adoption. Our work presents an active and contactless solution for drone sensing, positioning our system to readily benefit from ongoing advancements in event camera technology.